OpenStudy (anonymous):

I haven't done literal equations in a year so I totally forgot. I ALWAYS get overwhelmed by all of the things to do and by what order to do them in! Can someone help me? *My teacher also says that I have to factor binomials (and trinomials I'm guessing) so keep that in mind* xx

OpenStudy (anonymous):

Can you give me an example?

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

g=5x+xy+7y solve for x

OpenStudy (anonymous):

In terms of g and y?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok 1sec

OpenStudy (anonymous):

so bunch all the x's on one side. in this case, can be done by subtracting 7y to the left side, and you now have g-7y=5x+xy.

OpenStudy (anonymous):

now factor out x from the left side. now it looks like: g-7y=x(5+y)

OpenStudy (anonymous):

divide by (5+y) and you have (g-7y)/(5+y)=x

OpenStudy (anonymous):

That answer your question?

OpenStudy (anonymous):

Yup! I get it! Can you solve this example too? 2y - 2/3x + 6 = 0 solve for y

OpenStudy (anonymous):

can you put in the () where they go, b/c that could either be (2y-2)/(3x+6)=0 or 2y-(2/3)x+6=0

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

OK. You try it first, or at least tell me where you're getting stuck. I feel like I'm giving you all the answers, and I want you to learn something not just copy all the answers down

OpenStudy (anonymous):

Actually I'm doing them and then comparing your work to mine xD But cool here's my work: So I start with 2y-2/3x+6=0. I move over the 6 to the other side. 2y-2/3x=-6. I add -2/3x on both side to move it. 2y=-6+2/3x. I divide everything by 2. y=-3+1/3x. I'm really not sure how to divide a fraction so I just thought. What's half of 2/3? 1/3! Is that right? here's my final answer: y=-3+1/3x

OpenStudy (anonymous):

Yes, it is. Dividing fractions is very easy though, and you should know how to do it.

OpenStudy (anonymous):

1/2 * 1/2 is what? 1/2 divided by 1/2 is what?

OpenStudy (anonymous):

I know it's like multiply by negative reciprical but I don't get it very much :/

OpenStudy (anonymous):

not negative reciprocal just reciprocal :)

OpenStudy (anonymous):

1/2 by 1/2 is 1/4. 1/2 divided by 1/2 is 1?

OpenStudy (anonymous):

yes. so what's 45/79 divided by 11/37?

OpenStudy (anonymous):

and while you're at it, what's 3 and 4/5 divided by 5/6?

OpenStudy (anonymous):

Holy crap. Um, 1665/869? Do you mean 3 divided by 5/6 and 4/5 divided by 5/6? or 19/5 divided by 5/6?

OpenStudy (anonymous):

the 2nd one.

OpenStudy (anonymous):

114/25?

OpenStudy (anonymous):

So I just cross-multiply basically?

OpenStudy (anonymous):

it's a lot easier if you just flip the fraction you're dividing by and multiply.

OpenStudy (anonymous):

But yes, essentially the same thing.

OpenStudy (anonymous):

so do that in the previous problem 2/3 divided by 2.

OpenStudy (anonymous):

2/3 times 1/2 is 2/6. Simplified is 1/3. Okay! Well I got it right with my reasoning anyway but thanks for that! I really needed to know that(:

OpenStudy (anonymous):

great! i'm glad i could help.

OpenStudy (anonymous):

So did I get that first problem right? The one I explained step by step?

OpenStudy (anonymous):

Yes. You should have more confidence in your math skills. once you learn the basic equations and formulas, it's really easy.

OpenStudy (anonymous):

Ahaha thanks(: I just have a quiz tomorrow and I have the WORST memory. Let me see if I have this right- point slope form is y-y1=m(x-x1). Standard form is Ax+By=C *no fractions, x must be positive*. Direct variation y=kx k=y/x (use proportions). Inverse variation y=k/x k=yx. Literal equations factor binomials, to divide fracions multiply by reciprocal. AND THERE'S MY WHOLE NIGHT(:

OpenStudy (anonymous):

yeah :). keep in mind for the variation stuff all you really need to memorize is y=kx and k=yx for direct and inverse variation, respectively. from there you can divide by x by direct var. and get k=y/x or divide by x by inverse var. and get y=k/x.

OpenStudy (anonymous):

So basically dir. var. is y=kx. In. var. is y=k/x right?

OpenStudy (anonymous):

or k=yx. same thing.

OpenStudy (anonymous):

for which inverse variation?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

k=yx divide by x k/x=y Same thing!

OpenStudy (anonymous):

AND I'm done studying for the night.(: I'll be back soon ahaha

OpenStudy (anonymous):

Thank you SOOOOO much you have no idea what this means to me (:

OpenStudy (anonymous):

no problem. happy to help!